OpenStudy (anonymous):

Solve 2x + 3y = 18 for y.

OpenStudy (nubeer):

1. subtract 2x on both sides then 2. divide both sides by 3

OpenStudy (anonymous):

18-2x/3?

OpenStudy (nubeer):

yes that 3 would be in whole denominator i mean also below 18 (18-2x)/ 3

OpenStudy (anonymous):

y = 2x + 6 y = –2x – 6 y =-2/3 x + 6 y = 3/2x + 6 these are my answer choise, i've already done it the way you said but I cannot not get them answer they are looking for I was also thinking that the answer is the first one

OpenStudy (nubeer):

its tthe 3rd option.

OpenStudy (anonymous):

thanks man :)